OpenStudy (anonymous):

simplify sqrt 12+ sqrt 16 + sqrt 8 - sqrt 2...help meee....

OpenStudy (anonymous):

\[\sqrt{12}=2\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{16}=4\]

OpenStudy (anonymous):

\[\sqrt{8}=2\sqrt{2}\]

OpenStudy (anonymous):

\[4+\sqrt{2}+2\sqrt{3}\]

OpenStudy (anonymous):

\[2\sqrt{3}+4+2\sqrt{2}-\sqrt{2}\]

hartnn (hartnn):

did u understand how \(\sqrt {12} = 2 \sqrt 3\quad ??\)

OpenStudy (anonymous):

yes...thank you guys and hartnn it's sqrt 2* sqrt2 *sqrt 3

hartnn (hartnn):

yup, correct. u are smart :)